OpenStudy (anonymous):

please help me Radical 6 (5 radical 2+ radical 5)

OpenStudy (anonymous):

\[\sqrt{6}(5\sqrt{2}+\sqrt{5})\] this is what it looks like on the paper

OpenStudy (lynfran):

what do u have to do to this?

OpenStudy (anonymous):

multiply the radicals

OpenStudy (lynfran):

ok

OpenStudy (lynfran):

well we distribute \[\sqrt{6}*5\sqrt{2}+\sqrt{6}*\sqrt{5}\]

OpenStudy (anonymous):

ok then what do we do

OpenStudy (lynfran):

\[5\sqrt{2}*\sqrt6+\sqrt{6*5}\]\[5\sqrt{6*2}+\sqrt{30}\]

OpenStudy (lynfran):

\[5\sqrt{12}+\sqrt{30}\]\[5\sqrt{4*3}+\sqrt{30}\]\[10\sqrt{3}+\sqrt{30}\]

OpenStudy (anonymous):

and is that going to be the answer

OpenStudy (lynfran):

yes correct

OpenStudy (anonymous):

ok thank you very much

OpenStudy (lynfran):

welcome